OpenStudy (anonymous):

simplify: sqrt9x-8=sqrt10

OpenStudy (anonymous):

\[\sqrt{9x-8}=\sqrt{10}\]

OpenStudy (anonymous):

x = 1 x = 2 x = 12 None of the above i got d is that right

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Radicals on both sides? Godsend. Square both sides...

OpenStudy (jhannybean):

Yes @millikengal

OpenStudy (anonymous):

that was my second choice was b

OpenStudy (jhannybean):

Are you just guessing these or working them out? :P

OpenStudy (anonymous):

working someout guessing on the ones im not 100 percent sure about

OpenStudy (jhannybean):

ok :)